# MCP-Zero: Proactive Toolchain Construction for LLM Agents from Scratch

### v2

@@ 4.1 Data Collection and Scope
4.2 Data Processing and Quality Control
We implemented a rigorous filtering process to ensure data quality and completeness. Servers were excluded if they: (1) lacked MCP-compliant tool definitions, (2) missing documentation, or (3) failed to provide sufficient semantic information for retrieval purposes. This filtering process reduced our dataset from 396 to 308 high-quality servers.
For structured information extraction, we employed Qwen2.5-72B-Instruct [24] with manually crafted few-shot in-context learning examples to guide the extraction process. The extraction follows a standardized JSON schema:
@@ 4.2 Data Processing and Quality Control
A critical component of the MCP-tools dataset is the server_summary field, which we generate by leveraging LLMs to synthesize comprehensive summaries from README content. These summaries capture server capabilities, features, and usage patterns while excluding operational details like installation instructions. This design enables effective matching regardless of whether the model expresses tool requirements in formal descriptions or more specific, invocation-oriented language, thereby improving retrieval precision across diverse query formulations.
4.3 Dataset Statistics and Characteristics
The final MCP-tools dataset comprises 308 servers and 2,797 tools, representing a comprehensive coverage of the MCP ecosystem. The tool distribution exhibits significant variance: the mean number of tools per server is 9.08 (median: 5.0, standard deviation: 11.40). Over half of the servers (162 servers) contain 5 or fewer tools, while some specialized servers include over 60 tools, reflecting the diverse nature of MCP applications ranging from simple utilities to comprehensive API suites.
To support efficient retrieval, we pre-compute embeddings for all textual descriptions and summaries using OpenAI text-embedding-3-large, creating a searchable vector index that enables rapid semantic matching during the hierarchical routing process. These pre-computed embeddings are included as part of the dataset release.
To our knowledge, this represents the first dataset specifically designed for semantic tool retrieval in the MCP domain, complementing existing evaluation frameworks like MCPBench [13] by focusing on discovery rather than performance testing. The dataset’s design directly supports our MCP-Zero framework while remaining general enough to benefit other retrieval-based tool selection approaches. By providing both original MCP documentation and enhanced semantic summaries, MCP-tools enables researchers to explore various retrieval strategies and evaluate their effectiveness across diverse tool domains.
@@ 5.1 Existing Datasets and Their Limitations
5.2 Needle-in-a-Haystack Evaluation
  • MCP-Zero: Our proactive retrieval approach
@@ 5.3 APIBank Evaluation
  • Necessity of Proactive Requests. Pure query-retrieval baselines stall at 65–72 % accuracy, confirming that letting the model author semantically aligned requests is crucial.
Experiments on APIBank corroborate our claims: MCP-Zero delivers near-optimal or superior tool-selection accuracy while slashing context usage by up to two orders of magnitude, remaining robust in both single- and multi-turn conversations and under massive tool-pool scaling. These results highlight proactive, iterative tool discovery as a practical path toward scalable, cost-efficient agent systems.
@@ 6.1 Cookbook: Integrate MCP-Zero Into Agent
MCP-Zero is fundamentally a simple yet effective approach that we hope will benefit the broader MCP community. The core methodology distills into three straightforward steps: prompting models to proactively request tools, maintaining a lightweight tool index with semantic descriptions, and leveraging the improved semantic alignment for high-precision retrieval. Below we provide a practical guide for integrating these ideas into existing agent frameworks.
@@
In addition, the output structure needs to be specified as we mentioned in Section 3.1. This step aims to stimulate the model’s ability to ”proactively” propose requirements.
@@ 6.3 Synergy with Alita: Using and Making Tools
MCP-Zero and Alita address complementary halves of the same problem: the former maximises tool discovery while the latter maximises tool creation. When combined, they form a virtuous loop: an agent first proactively discover tools from all available resources; if none fits, it switches to Alita’s workflow to synthesize a new one, then registers the freshly built tool for the community. We believe such a pipeline is a compelling direction toward self-evolving, cost-aware agentic AI systems.
@@ 6.4 Future Work
MCP Server Implementation. A natural extension involves packaging MCP-Zero as a dedicated MCP server providing tool discovery services. This ”meta-server” would expose standardized APIs for proactive tool retrieval, enabling seamless integration into existing MCP ecosystems and serving as a centralized discovery hub for distributed tool collections.
Multi-Agent Orchestration. MCP-Zero’s proactive discovery approach could enable better multi-agent collaboration. Future work could investigate how different agents can automatically discover and share tools with each other, allowing them to work together more effectively on complex tasks that require diverse capabilities.

### v3

@@ 4.1 Dataset Construction
Structured Information Extraction. We employed Qwen2.5-72B-Instruct [24] with carefully designed few-shot examples to extract structured information following a standardized schema:
@@
The server_summary field represents a key innovation: LLM-generated comprehensive summaries that distill server capabilities while excluding operational details. This design enables effective semantic matching across diverse query formulations, from formal API specifications to natural language capability descriptions.
4.2 Dataset Characteristics
After the above process, we have completed the construction of the dataset as follows:
Scale and Diversity. The final dataset comprises 308 servers and 2,797 tools, providing comprehensive coverage of the MCP ecosystem. Tool distribution shows significant variance (mean: 9.08, median: 5.0, $\sigma$: 11.40), reflecting the ecosystem’s diversity from lightweight utilities to comprehensive API suites. Over half the servers (162) contain $\leq$5 tools, while specialized servers include 60+ tools.
Semantic Infrastructure. To support efficient retrieval evaluation, we pre-compute embeddings for all textual content using OpenAI text-embedding-3-large, creating a searchable vector index that enables rapid semantic matching. These embeddings are included in the dataset release to ensure reproducible evaluation.
Research Impact. MCP-tools represents the first dataset specifically designed for semantic tool discovery in the MCP domain, complementing performance-focused frameworks like MCPBench [13]. The dataset’s design supports not only our active discovery framework, but provides a general foundation for evaluating retrieval-based tool selection approaches across the research community.
@@ 5.1 Existing Datasets and Their Limitations
5.2 Needle-in-a-Haystack Evaluation
  • MCP-Zero: Our active retrieval approach
@@ 5.3 APIBank Evaluation
  • Necessity of Active Requests. Pure query-retrieval baselines stall at 65–72 % accuracy, confirming that letting the model author semantically aligned requests is crucial.
Experiments on APIBank corroborate our claims: MCP-Zero delivers near-optimal or superior tool-selection accuracy while slashing context usage by up to two orders of magnitude, remaining robust in both single- and multi-turn conversations and under massive tool-pool scaling. These results highlight active, iterative tool discovery as a practical path toward scalable, cost-efficient agent systems.
6 Conclusion
This work establishes active tool discovery as a fundamental paradigm for autonomous agent systems, enabling models to maintain decision autonomy while addressing critical scalability challenges in tool-calling architectures. MCP-Zero demonstrates that shifting from passive tool consumption to agent-driven capability acquisition achieves substantial efficiency gains—98% token reduction with preserved accuracy—while restoring the core principle of autonomous agency: the ability to assess limitations and actively acquire necessary resources. Our theoretical framework, empirical validation, and the MCP-tools dataset provide both the foundation and infrastructure for advancing autonomous agent architectures as tool ecosystems continue expanding exponentially.
@@ 7.1 Cookbook: Integrate MCP-Zero Into Agent
MCP-Zero is fundamentally a simple yet effective approach that we hope will benefit the broader MCP community. The core methodology distills into three straightforward steps: prompting models to actively request tools, maintaining a lightweight tool index with semantic descriptions, and leveraging the improved semantic alignment for high-precision retrieval. Below we provide a practical guide for integrating these ideas into existing agent frameworks.
@@
In addition, the output structure needs to be specified as we mentioned in Section 3.1. This step aims to stimulate the model’s ability to ”actively” propose requirements.
@@ 7.3 Synergy with Alita: Using and Making Tools
MCP-Zero and Alita address complementary halves of the same problem: the former maximises tool discovery while the latter maximises tool creation. When combined, they form a virtuous loop: an agent first actively discover tools from all available resources; if none fits, it switches to Alita’s workflow to synthesize a new one, then registers the freshly built tool for the community. We believe such a pipeline is a compelling direction toward self-evolving, cost-aware agentic AI systems.
@@ 7.4 Future Work
MCP Server Implementation. A natural extension involves packaging MCP-Zero as a dedicated MCP server providing tool discovery services. This ”meta-server” would expose standardized APIs for active tool retrieval, enabling seamless integration into existing MCP ecosystems and serving as a centralized discovery hub for distributed tool collections.
Multi-Agent Orchestration. MCP-Zero’s active discovery approach could enable better multi-agent collaboration. Future work could investigate how different agents can automatically discover and share tools with each other, allowing them to work together more effectively on complex tasks that require diverse capabilities.